The sum of two consecutive integers is 145. Find the two integers.
Answer by funmath(2933)

The sum of two consecutive integers is 145. Find the two integers.
Let the first number be: x
Then the second number is: x+1
Sum means: +
is means: =
(x)+(x+1)=145
2x+1=145
2x+1-1=145-1
2x=144
2x/2=144/2
x=72
The first number: x=72
The second number: x+1=72+1=73
:
Sanity check:
Are 72 and 73 consecutive integers? Yes!
Do they add up to 145? 72+73=145 Yes!
We don't appear to be insane.
Happy Calculating!!!